OpenStudy (abmon98):

When deriving the equations of motion in the x and y direction the resultant in the x direction is given by Rx=mx^..=mdx^./dx*dx/dt is the second equation derived from chain rule?

OpenStudy (irishboy123):

you might be veering toward this important result \(\ddot x = \dfrac{dv}{dt} = \dfrac{dv}{dx}.\dfrac{dx}{dt} = \dfrac{dv}{dx}.v = \dot x\dfrac{d \dot x}{dx}\) then again, you might not !! post / scan the thing as that might help :p